# Convergence of the extended Kalman filter with small and state-dependent noise

### v1

@@ 5 Impact of the initial error on the EKF estimation error
$|\pazocal{X}(t)|_{q}$$\leq$ $\sqrt{\varepsilon}C_{q}(\operatorname{tr}(\pazocal{K}(t)))^{1/2}$
$\leq$ $\sqrt{\varepsilon}C_{q}\left(1+q_{0}e^{(-(2c-k)t)}\right)^{1/2}$
$\leq$ $C_{q}\sqrt{\varepsilon}+C_{q}\sqrt{\varepsilon}|Y(0)-M(0)|_{q}e^{(-(c-k/2)t)}$
$\leq$ $C_{q}\sqrt{\varepsilon}+C_{q}|Y(0)-M(0)|_{q}e^{(-c_{0}t)},$
with $c_{0}=c-k/2$ and $\varepsilon$ small enough. ∎
Acknowledgements.
The authors thank the collaborators within the DFG research project “CESMO - Contain Epidemics with Stochastic Mixed-Integer Optimal Control”, in particular Gerd Wachsmuth, Armin Fügenschuh, Markus Friedemann, Jesse Beisegel (BTU Cottbus–Senftenberg), for insightful discussions and valuable suggestions that improved this paper. Funding  The authors gratefully acknowledge the support by the Deutsche Forschungsgemeinschaft (DFG), award number 458468407, and by the German Academic Exchange Service (DAAD), award number 57417894.
@@ Appendix B General results on symmetric positive definite (semi-definite) matrices
We refer to the book horn2012matrix for further details and proof of results in this section. Let $n,m$ be two positive integers, and $\pazocal{M}_{n}(\mathbb{R})$ the class of $n\times n$ matrices with real elements:
Definition B.1
@@ Appendix C The Loewner partial order
Proposition C.1
If $A\in\pazocal{M}_{n}(\mathbb{R})$ is symmetric with the smallest and largest eigenvalues $\lambda_{min}(A)$ and $\lambda_{max}(A)$, respectively, then $\lambda_{max}(A)I\succeq A\succeq\lambda_{min}(A)$
Proposition C.2
@@ Appendix D Some properties of the trace
We refer to gabih2020asymptotic for the proof of the following lemma.
Lemma D.1
@@
    $\lambda_{min}(A)\operatorname{tr}(B)\leq\operatorname{tr}(AB)\leq\lambda_{max}(A)\operatorname{tr}(B)$ (71)
@@
    $\dfrac{\operatorname{tr}(B)}{\operatorname{tr}\left(A^{-1}\right)}\leq\operatorname{tr}(AB)\leq\operatorname{tr}(A)\operatorname{tr}(B)$ (72)
@@
    $\operatorname{tr}^{2}(A)\geq\operatorname{tr}\left(A^{2}\right)\geq\dfrac{1}{d}\operatorname{tr}^{2}(A)$ (73)
@@
    $|A|_{F}=\sqrt{\operatorname{tr}\left(A^{2}\right)}\leq\operatorname{tr}(A)$ (74)
@@
$\operatorname{tr}(AB)\leq|\operatorname{tr}(AB)|\leq|A|_{2}\operatorname{tr}(B)$ (75)

### v2

@@ 5 Impact of the initial error on the EKF estimation error
$|\pazocal{X}(t)|_{q}~~\leq\sqrt{\varepsilon}C_{q}(\operatorname{tr}(\pazocal{K}(t)))^{1/2}$ $\leq\sqrt{\varepsilon}C_{q}\left(1+q_{0}e^{-(2c-k)t}\right)^{1/2}$
$\leq C_{q}\sqrt{\varepsilon}+C_{q}\sqrt{\varepsilon}|Y(0)-M(0)|_{q}e^{-(c-k/2)t}$
$\leq C_{q}\sqrt{\varepsilon}+C_{q}|Y(0)-M(0)|_{q}e^{-c_{0}t},$
with $c_{0}=c-k/2$ and $\varepsilon$ small enough.
Acknowledgments
The authors thank the collaborators within the DFG research project “CESMO - Contain Epidemics with Stochastic Mixed-Integer Optimal Control”, in particular Gerd Wachsmuth, Armin Fügenschuh, Markus Friedemann, Jesse Beisegel (BTU Cottbus–Senftenberg), for insightful discussions and valuable suggestions that improved this paper.
Funding  The authors gratefully acknowledge the support by the Deutsche Forschungsgemeinschaft (DFG), award number 458468407, and by the German Academic Exchange Service (DAAD), award number 57417894.
@@ Appendix B General results on symmetric positive definite (semi-definite) matrices
We refer to the book [14] for further details and proof of results in this section. Let $n,m$ be two positive integers, and $\pazocal{M}_{n}(\mathbb{R})$ the class of $n\times n$ matrices with real elements:
Definition B.1.
@@ Appendix C The Loewner partial order
Proposition C.1.
If $A\in\pazocal{M}_{n}(\mathbb{R})$ is symmetric with the smallest and largest eigenvalues $\lambda_{min}(A)$ and $\lambda_{max}(A)$, respectively, then $\lambda_{max}(A)I\succeq A\succeq\lambda_{min}(A)I$
Proposition C.2.
@@ Appendix D Some properties of the trace
We refer to [15] for the proof of the following lemma.
Lemma D.1.
@@
    $\lambda_{min}(A)\operatorname{tr}(B)\leq\operatorname{tr}(AB)\leq\lambda_{max}(A)\operatorname{tr}(B)$ (D.1)
@@
    $\dfrac{\operatorname{tr}(B)}{\operatorname{tr}\left(A^{-1}\right)}\leq\operatorname{tr}(AB)\leq\operatorname{tr}(A)\operatorname{tr}(B)$ (D.2)
@@
    $\operatorname{tr}^{2}(A)\geq\operatorname{tr}\left(A^{2}\right)\geq\dfrac{1}{d}\operatorname{tr}^{2}(A)$ (D.3)
@@
    $|A|_{F}=\sqrt{\operatorname{tr}\left(A^{2}\right)}\leq\operatorname{tr}(A)$ (D.4)
@@
$\operatorname{tr}(AB)\leq|\operatorname{tr}(AB)|\leq|A|_{2}\operatorname{tr}(B)$ (D.5)